Three identical small Styrofoam balls (m = 1.94g) are suspended from a fixed point by three nonconducting threads, each with a length of 54.7cm and with negligible mass. At equilibrium the three balls form an equilateral triangle with sides of 29.9cm. What is the common charge q carried by each ball?
1
Expert's answer
2020-01-20T05:24:17-0500


For the equilibrium on the y-axis:


"T\cos \gamma=mg"

For the equilibrium on the x-axis:

"T\sin \gamma=2F\cos {\frac{\alpha}{2}}=2\frac{kq^2}{a^2}\cos 30"

"\tan\gamma=2\frac{kq^2}{mga^2}\cos 30"

"\cos \gamma=\frac{b}{L}=\frac{\sqrt{3}a}{4L}=\frac{\sqrt{3}(0.299)}{4(0.547)}"

"\gamma=76.31\degree"

"\tan(76.31\degree)=2\frac{(8.99\cdot 10^9)q^2}{(0.00194)(9.8)(0.299)^2}\cos 30"

"q=1.59\cdot 10^{-6}C"